# Defining A New Cross Reality: Digital Twins and Mixed Reality Worlds

### v1

@@ CONCLUSION
This paper introduced the concept of cross-reality lifestyles and presented a framework for analyzing the integration of physical and virtual spaces in metaverse environments. The analysis reveals three distinct patterns of integration: amplification, complementary, and emergence, and identifies the technical requirements in the dimensions of content, platform infrastructure, and device interface. A systematic analysis of commercial metaverse platforms reveals how different platform types implement these requirements and support different forms of integration.
In the future, the boundaries between platforms are expected to become increasingly fluid and new hybrid forms will emerge, such as the fusion of industrial and social functions and the integration of gaming elements and practical applications. Future research should focus on developing metrics for evaluating cross-reality integration, investigating social and psychological effects, and exploring new interaction paradigms. We hope that the framework of this research will provide guidelines for platform selection and design of technical specifications for future metaverse services and application development.

### v2

@@ CONCLUSION
This paper introduced cross-reality lifestyles and the ACE Cube framework analyzing physical-virtual space integration. We introduced the "verse" concept, treating both spaces as activity spaces with different implementations, and analyzed integration patterns as continuous characteristics across Amplification, Complementary, and Emergence dimensions. The case analysis mapped experiences and technical requirements within the ACE space, establishing connections between integration patterns and implementation. Platform analysis revealed strategic differentiation along the A-C axis, with social VR platforms showing superior emergent capabilities. The ACE Cube framework provides guidelines for selecting platforms, prioritizing technical investments, and developing applications. Developers can identify technical requirements for target domains and formulate strategies for the ultimate metaverse platform.
Future research should focus on developing objective evaluation methods, exploring new interaction paradigms, and integrating high-ACE domains through emerging technologies. We expect this framework to guide the technical design of future metaverse services.